If the fuel consumption of a car, travelling at a steady 60 mph is 20 mpg, what is the power produced by the petrol being burnt in the engine? [Assume that the density of petrol is 0.84 kg/l and the heat of combustion of petrol is 47 302 kJ/kg.
1
Expert's answer
2019-01-04T10:51:10-0500

60 mph=26.67 m/s.

20 mpg=8443.27 m/L.

The volume in 1 second:

V=26.67/8443.27=0.0032 L.

The power produced by the petrol being burnt in the engine:

P=0.0032(0.84)(47302 )=125.5 kW.